The given statements relate to variables in the equation of exchange. Place the appropriate variable or variables next to each statement according to how a monetarist would describe the variable or variables money supply real GDP nominal GDP The average number of times MI Answer Bank each dollar is used in a vear total spending price level stable over time If M increases, this variable is most likely to increase as well.
Answer #1

According to Quantity theory of money, the relationship between money supply, velocity, price level and transactions

M*V=P*Q

Money Supply = M

Real GDP = Nominal GDP*price level = Q*P, quantity adjusted to price level

Nominal GDP = Q multiplied by current year prices

The average number of times = V

Total spending = Q

Price level = P

Stable over time = V

With the increase in money supply the inflation or price level too increases = P